OpenStudy (anonymous):

What is the function rule represented by the following table? x y 0 2 -1 -1 -2 -4

OpenStudy (tkhunny):

There are infinitely many answers to this question. Just find one and explain it.

OpenStudy (anonymous):

y=x+ 2 y= 3x+ 2 y= -3x+ 2 y= 2x is it any of these

OpenStudy (tkhunny):

If those are your only choices, try them out! Substitute the given values for 'x' and see if you get the corresponding value for 'y'.

OpenStudy (anonymous):

ok